the sum of 2 number is 100. The larger number is 10 more than twice the smaller
Answer by Schaman_Dempster(26)

Let x and y be the two numbers whose sum is 100 such that x>y.
So, x + y = 100 ------------- (1)
Also, larger number is 10 more than twice the smaller.
Therefore, x = 2y + 10
100 - y = 2y + 10 {Using equation (1) }
90 = 3y
Dividing both the sides by 3, we get:
90/3 = y
or y = 30
Plugging back the value of y in equation (1), we get:
x = 100 - y
x = 100 -30
x = 70
Thus, the larger number is 70 and the smaller number is 30.